OpenStudy (anonymous):

Use the given information to find the exact value of the expression. sin α =21/29 , α lies in quadrant II, and cos β = 15/17, β lies in quadrant I Find sin (α - β). A.155/493 B.468/493 C.475/493 D.132/493

OpenStudy (michele_laino):

from your data we have: \[\cos \alpha<0, \sin \beta>0\] Now, we can write: \[\cos \alpha=-\sqrt{1-(\frac{ 21 }{ 29 })^{2}}=-\frac{ 20 }{ 29 }\] and: \[\sin \beta=\sqrt{1-(\frac{ 15 }{ 17 })^{2}}=\frac{ 8 }{ 17 }\] so: \[\sin(\alpha-\beta)=\sin \alpha \cos \beta-\cos \alpha \sin \beta=\] \[=\frac{ 21 }{ 29 }*\frac{ 15 }{ 17 }+\frac{ 20 }{ 29 }*\frac{ 8 }{ 17 }=\frac{ 475 }{ 493 }\]

OpenStudy (michele_laino):

@SkiTTleoooo47

OpenStudy (anonymous):

Thank you!!!!!! :)

OpenStudy (michele_laino):

thank you!